# CollabLLM: From Passive Responders to Active Collaborators

### v2

@@ 3.2 Optimization & Synthetic Datasets
BLEU $\uparrow$ #Tokens$(k)\downarrow$ ITR $\uparrow$ PR $\uparrow$ #Tokens$(k)\downarrow$ ITR $\uparrow$ ACC $\uparrow$ #Tokens$(k)\downarrow$ ITR $\uparrow$
@@
Fine-tuning CollabLLMs. CollabLLMs are based on Llama-3.1-8B (Llama Team, 2024) with LoRA finetuning (Hu et al., 2022). We train four model variants: 1) Offline models: SFT and Offline DPO are fine-tuned on pre-generated multiturn conversational datasets guided by Multiturn-aware Rewards (MR) (cf. Section 3.2). 2) Online models: PPO and Online DPO are further trained from the SFT and Offline DPO models, respectively. The model during online fine-tuning is involved in the collaborative simulation to compute MRs, which, in turn, dynamically adjust the model preference.
Baselines. We compare CollabLLMs against (1) the pretrained Llama-3.1-8B (Base), (2) the base model with proactive prompt engineering (Proactive Base), which encourages follow-up and clarification questions.
@@ 5.3 Model Generalization (Table 2)
We assess CollabLLM, trained with online DPO on BigCodeBench-Chat (the coding assistance task), on Abg-CoQA (Guo et al., 2021), a question-answering (QA) benchmark where questions are labeled as ambiguous or non-ambiguous (cf. Appendix E). We categorize the model’s responses into two actions—asking a clarifying question or providing a direct answer—and evaluate action-level accuracy within each question type. As shown in Table 2, GPT-4o and Llama-3.1-8B rarely ask clarifying questions regardless of ambiguity. In contrast, CollabLLM proactively asks questions about 50% of the time while maintaining high accuracy on unambiguous inputs. This behavior leads to the highest Macro Accuracy across both ambiguous and non-ambiguous sets and improves Macro F1 over the base model, while leaving room for further improvement against GPT-4o. These results suggest that CollabLLM effectively generalizes its learned collaborative strategies beyond its training domain.
@@ 6 Real-world User Study
Quantitative Results (Figure 7). Across multiple metrics, CollabLLM consistently outperforms the baselines. It achieves an average document quality score of 8.50. Specifically, 91.4% of participants rate CollabLLM’s document quality as “good” (score 8–9), and 56.9% as “very good” (score 9–10), compared to 88.5% and 39.3% for Base (Llama-3.1-8B), respectively. Similarly, 63.8% of participants find CollabLLM highly engaging, while only 42.6% report the same for Llama-3.1-8B.
@@ Acknowledgments
We thank Doug Burger, Vishal Chowder, Jeevana Priya Inala, Hoifung Poon, Swadheen Shukla, Chandan Singh, Desney Tan and Chenglong Wang, as well as members of the Deep Learning and Health Futures groups at Microsoft Research for helpful discussions. We thank lab members in Leskovec and Zou’s labs for discussions and for providing feedback. We also gratefully acknowledge the support of NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM), IIS-2403318 (III); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
@@ Appendix C Safety Evaluation
As the models in this work are collaboratively trained and designed to be more aligned with the user’s intent, concerns may arise if a user happens to have malevolent intentions. However, we note that CollabLLM models were finetuned from Llama-3.1-8B, which has been aligned for safety—so jailbreaking CollabLLM still poses a significant challenge. To determine whether collaborative training weakens the safety features inherent to a model (Llama-3.1-8B) that has undergone significant alignment steps for safety, we performed an adversarial evaluation using the Azure AI Evaluation SDKhttps://learn.microsoft.com/en-us/python/api/overview/azure/ai-evaluation-readme and prompted both the baseline and CollabLLM with various offensive queries intended to elicit unsafe responses.
@@
  • Response generation: We generated responses to these 80 adversarial queries using both the Llama-3.1-8B baseline model and CollabLLM.
@@
Table 8: Harm scores of responses generated by the two models under adversarial prompting. Scores range from 00 to $7$, with values between 00 and $1$ indicating “very low” harm.
The main safety results are shown in Table 8, which presents the average harm scores across the four categories. Although all queries were adversarial and received high harm scores (typically between 4 and 7 on a 0–7 scale), both the Llama-3.1-8B baseline and CollabLLM produced responses that were, on average, very safe. Most scores are in the 0–1 range, which corresponds to “very low” harm. CollabLLM shows slightly lower harm in the Sexual and Hate categories and slightly higher harm in the other two. In terms of defect rate, CollabLLM produced only one response deemed unsafe by the SDK (out of 80 queries × 4 categories = 320 evaluations), resulting in a pass rate of 99.7%. Coincidentally, this is the same pass rate as Llama-3.1-8B, which also had one failed evaluation.

### v3

@@ 3.2 Optimization & Synthetic Datasets
[…]
@@
Fine-tuning CollabLLMs. CollabLLMs are based on Llama-3.1-8B-Instruct (Llama Team, 2024) with LoRA finetuning (Hu et al., 2022). We train four model variants: 1) Offline models: SFT and Offline DPO are fine-tuned on pre-generated multiturn conversational datasets guided by Multiturn-aware Rewards (MR) (cf. Section 3.2). 2) Online models: PPO and Online DPO are further trained from the SFT and Offline DPO models, respectively. The model during online fine-tuning is involved in the collaborative simulation to compute MRs, which, in turn, dynamically adjust the model preference.
Baselines. We compare CollabLLMs against (1) the pretrained Llama-3.1-8B-Instruct (Base), (2) the base model with proactive prompt engineering (Proactive Base), which encourages follow-up and clarification questions.
@@ 5.3 Model Generalization (Table 2)
We assess CollabLLM, trained with online DPO on BigCodeBench-Chat (the coding assistance task), on Abg-CoQA (Guo et al., 2021), a question-answering (QA) benchmark where questions are labeled as ambiguous or non-ambiguous (cf. Appendix E). We categorize the model’s responses into two actions—asking a clarifying question or providing a direct answer—and evaluate action-level accuracy within each question type. As shown in Table 2, GPT-4o and Llama-3.1-8B-Instruct rarely ask clarifying questions regardless of ambiguity. In contrast, CollabLLM asks questions about 50% of the time while maintaining high accuracy on unambiguous inputs. This behavior leads to the highest Macro Accuracy across both ambiguous and non-ambiguous sets and improves Macro F1 over the base model, while leaving room for further improvement against GPT-4o. These results suggest that CollabLLM effectively generalizes its learned collaborative strategies beyond its training domain.
@@ 6 Real-world User Study
Quantitative Results (Figure 7). Across multiple metrics, CollabLLM consistently outperforms the baselines. It achieves an average document quality score of 8.50. Specifically, 91.4% of participants rate CollabLLM’s document quality as “good” (score 8–9), and 56.9% as “very good” (score 9–10), compared to 88.5% and 39.3% for Base (Llama-3.1-8B-Instruct), respectively. Similarly, 63.8% of participants find CollabLLM highly engaging, while only 42.6% report the same for Base.
@@ Acknowledgments
We thank Doug Burger, Vishal Chowder, Jeevana Priya Inala, Giovanni Monea, Hoifung Poon, Swadheen Shukla, Chandan Singh, Alessandro Sordoni, Desney Tan and Chenglong Wang, as well as members of the Deep Learning and Health Futures groups at Microsoft Research for helpful discussions. We thank lab members in Leskovec and Zou’s labs for discussions and for providing feedback. We also gratefully acknowledge the support of NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM), IIS-2403318 (III); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
@@ Appendix C Safety Evaluation
As the models in this work are collaboratively trained and designed to be more aligned with the user’s intent, concerns may arise if a user happens to have malevolent intentions. However, we note that CollabLLM models were finetuned from Llama-3.1-8B-Instruct, which has been aligned for safety—so jailbreaking CollabLLM still poses a significant challenge. To determine whether collaborative training weakens the safety features inherent to a model (Llama-3.1-8B-Instruct) that has undergone significant alignment steps for safety, we performed an adversarial evaluation using the Azure AI Evaluation SDKhttps://learn.microsoft.com/en-us/python/api/overview/azure/ai-evaluation-readme and prompted both the baseline and CollabLLM with various offensive queries intended to elicit unsafe responses.
@@
  • Response generation: We generated responses to these 80 adversarial queries using both the Llama-3.1-8B-Instruct baseline model and CollabLLM.
@@
Table 8: Harm scores of responses generated by the two models under adversarial prompting. Scores range from 0 to 777, with values between 0 and 111 indicating “very low” harm.
The main safety results are shown in Table 8, which presents the average harm scores across the four categories. Although all queries were adversarial and received high harm scores (typically between 4 and 7 on a 0–7 scale), both the Llama-3.1-8B-Instruct baseline and CollabLLM produced responses that were, on average, very safe. Most scores are in the 0–1 range, which corresponds to “very low” harm. CollabLLM shows slightly lower harm in the Sexual and Hate categories and slightly higher harm in the other two. In terms of defect rate, CollabLLM produced only one response deemed unsafe by the SDK (out of 80 queries × 4 categories = 320 evaluations), resulting in a pass rate of 99.7%. Coincidentally, this is the same pass rate as Llama-3.1-8B-Instruct, which also had one failed evaluation.